# Structure-preserving space discretization of differential and nonlocal constitutive relations for port-Hamiltonian systems

### v1

@@ 2.2 Infinite-dimensional case
Difficulties arise when extending the pH framework to distributed pH systems. In particular, boundary terms and operator domains have to be taken into account. In this section, only a brief overview of the matter is given, the reader may refer to [46] for an extensive study along with well-posedness results. In particular, Stokes-Dirac structures [51] have been studied extensively as a special case of Dirac structures, where integration by parts (Stokes’ theorem) is encoded in the subspace, this allows the corresponding boundary terms to be expressed as boundary control port [14].
Recently, Stokes-Lagrange structures have been defined on $1$D domains [55] and $n$D domains [6] in order to adapt the Lagrange structure approach to infinite dimensional systems. Stokes-Lagrange structures are similar to Stokes-Dirac structures as they take into account the integration by parts. Moreover, they allow for the treatment of implicit constitutive relations in the distributed pH framework.
@@
Now, in order to define the Stokes-Dirac structure, we will consider three operators: $J:D(J)\subset\mathcal{E}_{s}\times\mathcal{E}_{R}\rightarrow\mathcal{E}_{s}\times\mathcal{E}_{R}$, along with observation/control operators $K:D(J)\rightarrow\mathcal{F}_{u}$ and $G:D(J)\rightarrow\mathcal{E}_{u}$ as defined in [6]. The property extensively used in the following and which should be emphasized here is the skew-symmetry:
$\forall z\in D(J),\quad\langle Jz,z\rangle_{\mathcal{E}_{s}}=\langle Gz,Kz\rangle_{\mathcal{E}_{u},\mathcal{F}_{u}}.$ (7)
@@
$\ker\begin{bmatrix}S_{|\mathcal{Z}_{0}}^{*}&-P_{|\mathcal{Z}_{0}}^{*}\end{bmatrix}\subset\textup{Ran}\begin{bmatrix}P\\ S\end{bmatrix}.$ (9)
@@
We can now state the definition of a distributed pH system:
@@
$\frac{\textup{d}}{\textup{d}t}H=\langle u_{D},y_{D}\rangle_{\mathcal{E}_{u},\mathcal{F}_{u}}+\langle\partial_{t}u_{L},y_{L}\rangle_{\mathcal{Y},\mathcal{U}}-\langle f_{R},Rf_{R}\rangle_{\mathcal{F}_{R},\mathcal{E}_{R}}.$ (12)
@@ 3.1 Nanorod equation
with $\varepsilon$ the strain, $p$ the linear momentum, $\sigma$ the stress, and $v$ the velocity. Now, we must define the constitutive relations that connect the state $(\varepsilon,p)$ and co-state $(\sigma,v)$ variables. In the linear local case, these are usually given by $v=p/\rho$ and Hooke’s law $\sigma=E\varepsilon$, where $E$ is the Young’s modulus and $\rho$ the mass density. Then, Eq. (13) becomes the classical wave equation. However, due to the scale of the studied system, these constitutive relations do not describe accurately the dynamics. To solve this issue, nonlocal effects between stress and strain variables have to be taken into account [30] in the constitutive relations.
@@ 3.2 Implicit Euler-Bernoulli beam
We focus now on the vibrations of a thin beam. However, it is well-known that the Euler-Bernoulli equation for this model is neither nonlocal nor implicit, see e.g. [26, 12]. Nevertheless, we recall that this is a matter of simplification in the limit of low wavenumbers, and that the shear model obtained by neglecting the inertia of cross section in the Timoshenko model is indeed implicit, as mentioned in [26, Eq. (13)]. It can also be derived in a more direct manner, as done now.
3.2.1 Modelling
Let us restart from the beginning, following [52, Chapter 2] to derive the model. Assume that the strain is proportional to the displacement, and that the linear filaments of the beam initially perpendicular to the middle line remain straight and perpendicular to the deformed middle line and undeformable: neither contraction nor extension (transverse shear effects are neglected).
Remark that the second assumption imposes a nonlinear relationship between the displacement $U(x,z,t)$, the displacement at a time $t$ of the particle which, when the beam is at equilibrium, occupies the position $(x,z)$, and $u(x,t)$, the displacement at a time $t$ occupying the position $(x,0)$ at equilibrium. If this relation is linearized, we obtain: $U=u-z\partial_{x}u_{z}.$
Denoting $w:=u_{z}$, the strain energy $\mathcal{P}$ of the beam is defined by:
where $L:=b-a$ is the length of the beam and $D:=E\frac{h^{3}}{12(1-\nu^{2})}$ is the modulus of flexural rigidity, with $E$ the Young’s modulus, $\nu$ the Poisson’s ratio ($0<\nu<\frac{1}{2}$ in physical situations), and $h$ the cross-sectional area ($h:=\pi r^{2}$ for cylindric beam of radius $r$). To include a tension $T_{0}$ as done in [9, Eq. (10)], and model pre-stressed beams, it requires adding more potential energy terms, following [26, Eq. (2)]. As this is not the purpose of this work, we restrict ourselves to the simpler models without tension.
The kinetic energy reads:
Principle of virtual work
$\mathcal{L}=\int_{0}^{T}\mathcal{K}+\mathcal{W}-\mathcal{P}\,\textup{d}t,$
Lemma 9.
Under the previously listed assumption, the equation of motion for the Euler-Bernoulli beam reads:
@@ 3.2.1 Modelling
3.2.2 Port-Hamiltonian representation
Let us now recast the implicit Euler-Bernoulli model as a pH system.
@@ 3.2.2 Port-Hamiltonian representation
$\begin{array}[]{c}(u_{L},y_{L})=\left(\begin{pmatrix}v(a)\\ v(b)\end{pmatrix},\frac{\rho h^{3}}{12}\,\begin{pmatrix}-\partial_{x}v(a)\\ \partial_{x}v(b)\end{pmatrix}\right),\\ (u^{1}_{D},y^{1}_{D}):=\left(\begin{pmatrix}\sigma(a)\\ \sigma(b)\end{pmatrix},\begin{pmatrix}-\partial_{x}v(a)\\ \partial_{x}v(b)\end{pmatrix}\right).(u^{2}_{D},y^{2}_{D}):=\left(\begin{pmatrix}v(a)\\ v(b)\end{pmatrix},\begin{pmatrix}\partial_{x}\sigma(a)\\ -\partial_{x}\sigma(b)\end{pmatrix}\right).\end{array}$ (28)
@@
This model was studied previously in, e.g.,  [25, 26] (with a tension $T_{0}\neq 0$). It was demonstrated that the behavior of the phase velocity and group velocity is closer to the Timoshenko beam than the classical Euler-Bernoulli beam model. Additionally, due to the non-trivial P operator, and considering a solution given as a monochromatic wave, $e^{i(kx-\omega t)}$, the phase velocity $\frac{\omega}{k}={k\sqrt{D}}/\sqrt{\rho h(1+(h^{2}/12)\,k^{2})}$ remains bounded at high frequency, which is not the case for the classical Euler-Bernoulli beam model. These properties advocate for the use of this model, especially for medium-sized beams.
@@ 3.3.1 Port-Hamiltonian representation
$\rho_{0}(\partial_{t}+\bm{u}\cdot\textbf{{grad}})\bm{u}&=-\textbf{{grad}}(p)+\mu\bm{\Delta u},\\ \operatorname{div}(u)&=0\end{aligned}\right.$ (29)
@@
In particular, $\textup{curl}_{2D}\textbf{{grad}}=0$, $\operatorname{div}\textbf{{grad}}^{\perp}=0$, moreover, $-\operatorname{div}$ is the formal adjoint of grad and $\textbf{{grad}}^{\perp}$ is the formal adjoint of $\textup{curl}_{2D}$.
@@
$\mathcal{K}(\psi):=\frac{\rho_{0}}{2}\int_{\Omega}||\textbf{{grad}}^{\perp}\psi||^{2}\,\textup{d}x,\quad\mathcal{E}(\omega):=\frac{\rho_{0}}{2}\int_{\Omega}\omega^{2}\,\textup{d}x.$ (31)
@@
$\frac{\textup{d}}{\textup{d}t}\mathcal{E}(\omega)=-\mu\int_{\Omega}||\textbf{{grad}}(\omega)||^{2}\,\textup{d}x-\int_{\partial\Omega}\underbrace{\rho_{0}\,\omega\,\psi}_{=:y_{D}^{4}}\,\underbrace{\bm{t}\cdot\textbf{{grad}}(\omega)}_{=:u_{D}^{4}}\,\textup{d}s\,-\int_{\partial\Omega}\underbrace{\omega}_{=:y_{D}^{5}}\,\underbrace{\mu\,\bm{n}\cdot\textbf{{grad}}(\omega)}_{:=u_{D}^{5}}\,\textup{d}s.$ (36)
In particular, with zero boundary conditions, and using the fact that $-\Delta\psi=\omega$, one gets:
$\frac{\textup{d}}{\textup{d}t}\mathcal{K}(\psi)=-2\,\mu\,\mathcal{E}(\omega),\qquad\frac{\textup{d}}{\textup{d}t}\mathcal{E}(\omega)=-\mu\int_{\Omega}||\textbf{{grad}}(\omega)||^{2}\,\textup{d}x.$
@@ 3.3.2 Boundary conditions
Using $-\Delta\psi=\omega,$ additional constraints are derived: $u_{D}^{3}=-\mu\,y_{D}^{5},\quad u_{D}^{5}=-u_{D}^{2}.$ Finally, gathering these conditions yields:
@@ 3.3.3 Enstrophy creation and kinetic energy dissipation
$\frac{\textup{d}}{\textup{d}t}\mathcal{E}(\omega)=-\mu\int_{\Omega}||\textbf{{grad}}(\omega)||^{2}\,\textup{d}x-\frac{1}{\mu}\int_{\partial\Omega}u_{D}^{5}u_{D}^{3}\,\textup{d}s.$ (38)
@@ 4.1 Nanorod equation
The structure-preserving spatial discretization of the wave equation as a pH system is already well known (see e.g., [41]). Here the additional difficulty is the implicit constitutive relation. In order to properly discretize (23) into a finite-dimensional pH system, we will take care of preserving the symmetry and skew-symmetry of the Lagrange structure and Dirac structure matrices, respectively. Moreover, one should be particularly cautious of the Robin boundary condition (17) as it will be of importance both in the Dirac and Lagrange structure discretizations. In particular, we will first write the weak formulation without these boundary conditions in order to identify the control ports at the discrete level, then apply these conditions.
@@ 4.1.1 Weak formulation
Carrying both integrations by parts yields:
$\begin{cases}\partial_{t}\int_{a}^{b}\,\frac{1}{E}(\,\phi\,\sigma+\ell^{2}\,\partial_{x}\phi\,\partial_{x}\sigma)\,\textup{d}x\,\underbrace{-[\phi\frac{\ell^{2}}{E}\,\partial_{x}\partial_{t}\sigma]_{a}^{b}}_{\text{Energy control port}}=\int_{a}^{b}\phi\,\partial_{x}v\,\textup{d}x,\\ \partial_{t}\int_{a}^{b}\phi\,\rho\,\partial_{t}v\,\textup{d}x=-\int_{a}^{b}\partial_{x}\phi\,\sigma\,\textup{d}x\,\underbrace{+\,[\phi\sigma]_{a}^{b}}_{\text{Power control port}}.\end{cases}$ (40)
@@
Before discretizing the system, let us finally apply the Robin boundary conditions. Denoting by $n_{a}=-1,\,n_{b}=1$ the outer normals, we get on the first and second lines:
@@
where the last equality is due to the equality $\rho\partial_{t}v=\partial_{x}\sigma$. Using these two results gives us:
$\begin{cases}\partial_{t}\int_{a}^{b}\,\frac{1}{E}(\,\phi\,\sigma+\ell^{2}\,\partial_{x}\phi\,\partial_{x}\sigma)\,\textup{d}x\,+\phi(a)\frac{\ell}{E}\,\partial_{t}\sigma(t,a)+\phi(b)\frac{\ell}{E}\,\partial_{t}\sigma(t,b)=\int_{a}^{b}\phi\,\partial_{x}v\,\textup{d}x,\\ \partial_{t}\int_{a}^{b}\phi\,\rho\,\partial_{t}v\,\textup{d}x+\phi(a)\,\ell\rho\partial_{t}v(t,a)+\phi(b)\,\ell\rho\partial_{t}v(t,b)=-\int_{a}^{b}\partial_{x}\phi\,\sigma\,\textup{d}x.\end{cases}$
Note that since the conditions are homogeneous, the controls do not appear in the equations anymore. For the sake of completeness and clarity, we will now carry the discretization without the Robin boundary conditions, then in a second stage, apply said conditions.

### v2

@@ 2.3 Infinite-dimensional case
Difficulties arise when extending the pH framework to distributed pH systems , i.e., when the system’s states also depend on a space variable. In particular, boundary terms and operator domains have to be taken into account. In this section, only a brief overview of the matter is given, the reader may refer to [46] for an extensive study along with well-posedness results. In particular, Stokes-Dirac structures [51] have been studied extensively as a special case of Dirac structures, where integration by parts (Stokes’ theorem) is encoded in the subspace, this allows the corresponding boundary terms to be expressed as boundary control port [14].
Recently, Stokes-Lagrange structures have been defined on 11D domains [55] and nnD domains [6] in order to adapt the Lagrange structure approach to infinite dimensional systems. Stokes-Lagrange structures are similar to Stokes-Dirac structures as they take into account the integration by parts. Moreover, they allow for the treatment of implicit constitutive relations in the distributed pH framework.
Definition 4.
A Stokes-Dirac structure (respectively, a Stokes-Lagrange structure) is a Dirac structure (respectively, a Lagrange structure) of infinite dimension that makes use of the Stokes’ divergence theorem (i.e., integration by parts).
@@
Now, in order to define the Stokes-Dirac structure, we will consider three operators: $J:D(J)\subset\mathcal{E}_{s}\times\mathcal{E}_{R}\rightarrow\mathcal{E}_{s}\times\mathcal{E}_{R}$, along with observation/control operators $K:D(J)\rightarrow\mathcal{F}_{u}$ and $G:D(J)\rightarrow\mathcal{E}_{u}$ as defined in [6] , i.e., such that (7) holds and $\begin{bmatrix}K\\ G\end{bmatrix}$ is onto in $\mathcal{F}_{u}\times\mathcal{E}_{u}$. The property extensively used in the following and which should be emphasized here is the skew-symmetry:
$\forall z\in D(J),\quad\langle Jz,z\rangle_{\mathcal{E}_{s}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\times\mathcal{E}_{R}}}=\langle Gz,Kz\rangle_{\mathcal{E}_{u},\mathcal{F}_{u}}.$ (7)
This latter identity, an Abstract Stokes’ divergence theorem, is indeed a skew-symmetry property if one considers it on the subspace $\mathcal{W}_{0}:=\ker K\cap\ker G$. Furthermore, if $J_{0}:=J|_{\mathcal{W}_{0}}$ satisfies $J_{0}^{*}=-J$ (the star denotes the adjoint operator), then $\mathcal{D}:=\begin{bmatrix}J\\ K\\  \operatorname{Id}\\ G\end{bmatrix}\mathcal{D}(J)$ is a Stokes-Dirac structure, see [51].
@@
$\ker\begin{bmatrix}S_{|\mathcal{Z}_{0}}^{*}&-P_{|\mathcal{Z}_{0}}^{*}\end{bmatrix}\subset\textup{Ran}\begin{bmatrix}P\\ S\end{bmatrix}$, (9)
where the star denotes the adjoint operator.
@@
We can now state the definition of a distributed pH system.
@@
Furthermore, we point out that Definition 5 is indeed an infinite-dimensional counterpart to Definition 1, but making use of the latent state variable to avoid the adjoint of $P$ and $S$, and with control and observation boundary terms expressed separately from $J$ (for $u_{D}$ and $y_{D}$) and $P$ and $S$ (for $u_{L}$ and $y_{L}$), as often in boundary control systems theory.
@@
$\frac{\textup{d}}{\textup{d}t}H=\langle u_{D},y_{D}\rangle_{\mathcal{E}_{u},\mathcal{F}_{u}}+\langle\partial_{t}u_{L},y_{L}\rangle_{\mathcal{Y},\mathcal{U}}-\langle f_{R},Rf_{R}\rangle_{\mathcal{F}_{R},\mathcal{E}_{R}}.$
@@
Remark 6.
It is important to notice that (11) is not the unique Hamiltonian that can be defined for system (10). Even if one imposes the Hamiltonian to be a quadratic functional, it remains non unique. Indeed, a Legendre transform on the boundary term $\gamma z$ provides the alternative definition:
$\widetilde{H}(z):=\frac{1}{2}\langle Pz,Sz\rangle_{\mathcal{X},\mathcal{E}_{s}}-\frac{1}{2}\langle\gamma z,\beta z\rangle_{\mathcal{Y},\mathcal{U}},$
for which the power balance becomes:
$\frac{\textup{d}}{\textup{d}t}\widetilde{H}=\langle u_{D},y_{D}\rangle_{\mathcal{E}_{u},\mathcal{F}_{u}}-\langle u_{L},\partial_{t}y_{L}\rangle_{\mathcal{Y},\mathcal{U}}-\langle f_{R},Rf_{R}\rangle_{\mathcal{F}_{R},\mathcal{E}_{R}}.$
From a control perspective, choosing $\widetilde{H}$ or $H$ gives us access to different input and output variables.
@@ 3.1 Nanorod equation
together with constitutive relations defined later, with $\varepsilon$ the strain, $p$ the linear momentum, $\sigma$ the stress, and $v$ the velocity. Now, we must define the constitutive relations that connect the state $(\varepsilon,p)$ and co-state $(\sigma,v)$ variables. In the linear local case, these are usually given by $v=p/\rho$ and Hooke’s law $\sigma=E\varepsilon$, where $E$ is the Young’s modulus and $\rho$ the mass density. Then, Eq. (12) becomes the classical wave equation. However, due to the scale of the studied system, these constitutive relations do not describe accurately the dynamics. To solve this issue, nonlocal effects between stress and strain variables have to be taken into account [30] in the constitutive relations.
@@ 3.2 Implicit Euler-Bernoulli beam
We focus now on the vibrations of a thin beam. However, it is well-known that the Euler-Bernoulli equation for this model is neither nonlocal nor implicit, see e.g. [26, 12]. Nevertheless, we recall that this is a matter of simplification in the limit of low wavenumbers, and that the shear model obtained by neglecting the inertia of cross section in the Timoshenko model is indeed implicit, as mentioned in [26, Eq. (13)]. It can also be derived in a more direct manner, as done in D.
Without external force, the equation of motion for the Euler-Bernoulli beam (54) reads:
@@
Let us recast it as a pH system.
@@
Proof.
Once the choice of energy variable has been made, the port-Hamiltonian formulation of the system is simply a reformulation of the original system. ∎
@@
Proof.
A direct computation yields the desired result. ∎
@@
[…]
@@
This model was studied previously in, e.g.,  [25, 26] (with a tension $T_{0}\neq 0$). It was demonstrated that the behavior of the phase velocity and group velocity is closer to the Timoshenko beam than the classical Euler-Bernoulli beam model. Additionally, due to the non-trivial P operator, and considering a solution given as a monochromatic wave, $e^{i(kx-\omega t)}$, the phase velocity $\frac{\omega}{k}={k\sqrt{D}}/\sqrt{\rho h(1+(h^{2}/12)\,k^{2})}$ remains bounded at high frequency, which is not the case for the classical Euler-Bernoulli beam model. These properties advocate for the use of this model, especially for beams of moderate thickness.
@@ 3.3.1 Port-Hamiltonian representation
$\left\{\begin{aligned} \rho_{0}(\partial_{t}+\bm{u}\cdot\textbf{{grad}})\bm{u}&=-\textbf{{grad}}(p)+\mu\bm{\Delta u},\\ \operatorname{div}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{u}})&=0\end{aligned}\right$. (28)
@@
In particular, $\textup{curl}_{2D}\textbf{{grad}}=0$, $\operatorname{div}\textbf{{grad}}^{\perp}=0$, moreover, $\operatorname{div}$ is the formal adjoint of grad and $\textbf{{grad}}^{\perp}$ is the formal adjoint of $\textup{curl}_{2D}$.
@@
[…]
@@
Proof.
First, writing (31) using $\psi$ instead of $\omega$ as the state variable gives us a co-energy formulation:
$\rho_{0}\,\Delta\partial_{t}\psi=-\rho_{0}\,\operatorname{div}(\omega\,\textbf{{grad}}^{\perp}(e_{\mathcal{K}}))-\mu\Delta^{2}\psi,\quad\text{with}\quad e_{\mathcal{K}}=\psi.$
Then, recalling that $J_{\omega}=\textbf{{curl}}_{2D}(G(\omega)\textbf{{grad}}^{\perp}(\cdot))=\operatorname{div}(\,\omega\,\textbf{{grad}}^{\perp}(\cdot))$, gives us the first line of (33).
Most notably, $\psi$ is the energy variable, the kinetic energy $\mathcal{K}$ is the Hamiltonian, and $\omega$ modulates the Dirac structure.
Then, the dynamics of the vorticity $\omega$ is given directly by (32), where the enstrophy $\mathcal{E}$ is the Hamiltonian, with $\omega$ as energy variable and $\psi$ modulates the Dirac structure.
@@
[…]
In particular, with zero boundary conditions, and using the fact that $\Delta\psi=\omega$, one gets:
[…]
Proof.
A direct computation yields the desired result. ∎
@@ 3.3.2 Boundary conditions
Using $\Delta\psi=\omega,$ additional constraints are derived: $u_{D}^{3}=-\mu\,y_{D}^{5},\quad u_{D}^{5}=-u_{D}^{2}.$ Finally, gathering these conditions yields:
@@ 3.3.3 Enstrophy creation and kinetic energy dissipation
[…]
@@
Proof.
Using the power balance provided in Theorem 14 with the boundary conditions of Lemma 15 yields the desired result. ∎
@@ 4.1 Nanorod equation
The structure-preserving spatial discretization of the wave equation as a pH system is already well known (see e.g., [41]). Here the additional difficulty is the implicit constitutive relation. In order to properly discretize (22) into a finite-dimensional pH system, we will take care of preserving the symmetry and skew-symmetry of the Lagrange structure and Dirac structure matrices, respectively. Moreover, one should be particularly cautious of the Robin boundary condition (16) as it will be of importance both in the Dirac and Lagrange structure discretizations. In particular, we will first write the weak formulation with general boundary conditions in order to exhibit the power control port and energy control port at the discrete level, then apply these conditions. As a consequence, one first retrieves the energy and power control ports that correspond to the continuous case, then closes the system using provided boundary conditions.
@@ 4.1.1 Weak formulation
Carrying out both integrations by parts yields:
$\begin{cases}\partial_{t}\int_{a}^{b}\,\frac{1}{E}(\,\phi\,\sigma+\ell^{2}\,\partial_{x}\phi\,\partial_{x}\sigma)\,\textup{d}x\,\underbrace{-[\phi\frac{\ell^{2}}{E}\,\partial_{x}\partial_{t}\sigma]_{a}^{b}}_{\text{Energy control port}}=\int_{a}^{b}\phi\,\partial_{x}v\,\textup{d}x,\\ \partial_{t}\int_{a}^{b}\phi\,\rho v\,\textup{d}x=-\int_{a}^{b}\partial_{x}\phi\,\sigma\,\textup{d}x\,\underbrace{+\,[\phi\sigma]_{a}^{b}}_{\text{Power control port}}.\end{cases}$ (39)
@@
Before discretizing the system, let us finally apply the Robin boundary conditions  (16). Denoting by $n_{a}=-1,\,n_{b}=1$ the outer normals, we get for the integrated terms of (39):
@@
where the latter equality is due to the equality $\rho\partial_{t}v=\partial_{x}\sigma$. Using these two results gives us:
$\begin{cases}\partial_{t}\int_{a}^{b}\,\frac{1}{E}(\,\phi\,\sigma+\ell^{2}\,\partial_{x}\phi\,\partial_{x}\sigma)\,\textup{d}x\,+\phi(a)\frac{\ell}{E}\,\partial_{t}\sigma(t,a)+\phi(b)\frac{\ell}{E}\,\partial_{t}\sigma(t,b)=\int_{a}^{b}\phi\,\partial_{x}v\,\textup{d}x,\\ \partial_{t}\int_{a}^{b}\phi\,\rho v\,\textup{d}x+\phi(a)\,\ell\rho v(t,a)+\phi(b)\,\ell\rho v(t,b)=-\int_{a}^{b}\partial_{x}\phi\,\sigma\,\textup{d}x.\end{cases}$
Note that since the conditions are homogeneous, the controls do not appear in the equations anymore. For the sake of completeness and clarity, we will now carry the discretization without the Robin boundary conditions, then in a second stage, apply these conditions.